# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ Abstract
We propose a method for simultaneously estimating a contemporaneous graph structure and autocorrelation structure for a causal high-dimensional vector autoregressive process (VAR). The graph is estimated by estimating the stationary precision matrix using a Bayesian framework. We introduce a novel parameterization that is convenient for jointly estimating the precision matrix and the autocovariance matrices. The methodology based on the new parameterization has several desirable properties. A key feature of the proposed methodology is that it maintains causality of the process in its estimates and also provides a fast feasible way for computing the reduced rank likelihood for a high-dimensional Gaussian VAR. We use sparse priors along with the likelihood under the new parameterization to obtain the posterior of the graphical parameters as well as that of the temporal parameters. An efficient Markov Chain Monte Carlo (MCMC) algorithm is developed for posterior computation. We also establish theoretical consistency properties for the high-dimensional posterior. The proposed methodology shows excellent performance in simulations and real data applications.
@@ 1 Introduction
Graphical models are popular models that encode scientific linkages between variables of interest through conditional independence structure and provide a parsimonious representation of the joint probability distribution of the variables, particularly when the number of variables is large. Thus, learning the graph structure underlying the joint probability distribution of a large collection of variables is an important problem and it has a long history. When the joint distribution is Gaussian, learning graphical structure can be done by estimating the precision matrix leading to the popular Gaussian Graphical Model (GGM); even in the non-Gaussian case one can think of encoding the graph using the partial correlation structure and learn the Partial Correlation Graphical Model (PCGM) by estimating the precision matrix. However, when the sample has temporal dependence, estimating the precision matrix efficiently from the sample can be challenging. If estimating the graphical structure is the main objective, the temporal dependence can be treated as a nuisance feature and ignored in the process of graph estimation. When the temporal features are also important to learn, estimating the graph structure and the temporal dependence simultaneously can be considerably more complex.
@@
There are several formulations of graphical models for time series. A time series graph could be one with the nodes representing the entire coordinate processes. In particular, if the coordinate processes are jointly stationary, this leads to a ‘stationary graphical’ structure. Conditional independence in such a graph can be expressed equivalently in terms of the absence of partial spectral coherence between two nodal series; see Dahlhaus (2000). Estimation of such stationary graphs have been investigated by Jung et al. (2015); Fiecas et al. (2019), Basu and Rao (2023), Basu et al. (2015) and Ma and Michailidis (2016).
@@
A weaker form of conditional coding for stationary multivariate time series is a ‘contemporaneous stationary graphical’ structure where the graphical structure is encoded in the marginal precision matrix. In a contemporaneous stationary graph, the nodes are the coordinate variables of the vector time series. If the time series is Gaussian, this leads to contemporaneous conditional independence among the coordinates of the vector time series. This is the GGM structure which has the appealing property that conditional independence of a pair of variables given the others is equivalent to the zero-value of the corresponding off-diagonal entry of the precision matrix and hence graph learning can be achieved via estimation of precision matrices. A rich literature on estimating the graphical structure can be found in classical books like Lauritzen (1996); Koller and Friedman (2009); see also Maathuis et al. (2019) and references therein for more recent results. When the distribution is specified only up to the second moment, which is common in the study of second-order stationary time series, one could learn the contemporaneous partial correlation graphical structure by estimating the stationary precision matrix. In this paper, we use the contemporaneous precision matrix of a stationary time series to estimate the graph along the line of Qiu et al. (2016); Zhang and Wu (2017). For modeling the temporal dynamics we use a vector autoregression (VAR) model.
@@
This paper brings together two popular models: the PCGM for graph estimation and the VAR model for estimating the temporal dynamics in vector time series. However, modeling the partial correlation graph and the temporal dependence with the VAR structure simultaneously is challenging, particularly under constraints such as reduced rank and causality on the VAR model and sparsity on the graphical model. This is achieved in this paper via a novel parameterization of the VAR process. The main contribution of the paper is a methodology that allows for meeting the following two challenging objectives simultaneously:
@@
    Estimation of the contemporaneous stationary graph structure under a sparsity constraint.
@@
    Estimation of VAR processes with a reduced rank structure under causality constraints.
@@
    establish posterior concentration under priors based on new parameterization.
@@ 2 Partial Correlation Graph Under Autoregression
In many contexts, the set of variables of interest evolve over time and are temporally dependent. If the process $\{\bm{X}_{t}:t=0,1,2,\ldots\}$ is stationary, that is, the joint distributions remain invariant under a time-shift, the relational graph of $\bm{X}_{t}$ remains time-invariant. A vector autoregressive (VAR) process provides a simple, interpretable mechanism for temporal dependence by representing the process as a fixed linear combination of itself at a few immediate time points plus an independent random error. It is widely used in time series modeling. In this paper, we propose a Bayesian method for learning the relational graph of a stationary VAR process.
@@
Causality is a property that plays an important role in multivariate time series models, particularly in terms of forecasting. For a causal time series, the prediction formula includes current and past innovations, and hence a causal time series allows a stable forecast of the future in terms of present and past data. However, the condition of causality imposes complex constraints on the parameters of the process, often making it extremely difficult for one to impose causality during the estimation process. An effective approach is to parameterize the constrained parameter space of causal processes in terms of unconstrained parameters and write the likelihood and the prior distribution in terms of the unconstrained parameters.
@@
$\det(\bm{A}(z))=0\mbox{ where }\bm{A}(z)=\bm{I}-\bm{A}_{1}z-\cdots-\bm{A}_{p}z^{p},\;z\in{\mathbb{C}}$ (2.2)
is the matrix polynomial associated with the VAR equation. For the VAR model to be causal, all roots of the determinantal equation must lie outside the unit disc $\mathcal{D}=\{z\in{\mathbb{C}}:|z|\leq 1\}$. When the roots of the determinantal equation lie outside the unit disc, the associated monic matrix polynomial ${\tilde{\bm{A}}}(z)=z^{p}-\bm{A}_{1}z^{p-1}-\cdots-\bm{A}_{p}$ is called Schur stable. Roy et al. (2019) provided a parameterization of the entire class of Schur stable polynomials and thereby parameterized the space of causal VAR models.
@@
Our main objectives in this paper is to estimate the stationary precision matrix of the process, $[{\rm Var}(X_{t})]^{-1}=\Omega=\bm{\Gamma}^{-1}(0)$, under sparsity restrictions. In the Roy et al. (2019) parameterization, the stationary variance $\bm{\Gamma}(0)$ and hence $\bm{\Omega}$ are functions of the basic free parameters and hence not suitable for estimation of the graphical structure with desired sparsity properties. We suggest a novel modification of the previous parameterization that achieves the goal of parameterizing the graphical structure and the VAR correlation structure directly and separately, thereby facilitating the estimation of both components under the desired restrictions, even in higher dimensions.
@@ 3 Reduced Rank Parameterization and Priors
The main idea in the proposed parameterization is separation of the parameters pertaining to the graphical structure, $\bm{\Omega},$ and the parameters that are used to describe the temporal correlation present in the sample. The following result is essential in developing the parameterization. It follows easily from (2.5) but due to its central nature in the new parameterization, we state it formally.
@@
Constraint-free parameterization: Based on Proposition 1, for a causal VAR process the successive differences $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ are nonnegative definite for $j=1,\ldots,p$, where $\bm{\Omega}=\bm{C}_{0}^{-1}$. Hence, these successive differences, along with the precision matrix $\bm{\Omega}$, can be parameterized using an unrestricted parameterization that maps non-negative definite matrices to free parameters.

### v2

# Relational Graph in Vector Autoregression: A Case Study on the Effect of the Great Recession on Connectivity of Economic Indicators

@@ Abstract
Under a high-dimensional vector autoregressive (VAR) model, we propose a way of efficiently estimating both the stationary graph structure between the nodal time series and their temporal dynamics. The framework is then used to make inferences on the change in interdependencies between several economic indicators due to the impact of the Great Recession, the financial crisis that lasted from 2007 through 2009. There are several key advantages of the proposed framework; (1) it develops a reparametrized VAR likelihood that can be used in general high-dimensional VAR problems, (2) it strictly maintains causality of the estimated process, making inference on stationary features more meaningful and (3) it is computationally efficient due to the reduced rank structure of the parameterization. We apply the methodology to the seasonally adjusted quarterly economic indicators available in the FRED-QD database of the Federal Reserve. The analysis essentially confirms much of the prevailing knowledge about the impact of the Great Recession on different economic indicators. At the same time, it provides deeper insight into the nature and extent of the impact on the interplay of the different indicators. We also contribute to the theory of Bayesian VAR by showing the consistency of the posterior under sparse priors for the parameters of the reduced rank formulation of the VAR process.
@@ 1 Introduction
Large dynamic systems like the US economy go through periods of upturns and downturns. However, extreme events such as the Great Recession and the COVID-19 epidemic cause upheavals beyond the reach of normal cycles. The Great Recession disrupted traditional and well-understood relationships among U.S. economic indicators, leading to unconventional dynamics in and among economic, labor, and financial markets during and after the event. Despite several government interventions, the sheer magnitude of the recession made recovery extremely hard. It was posited that many of the market indicators went through permanent structural changes as the economy emerged from the recession. The recession not only had a disparate impact on different market sectors but also restructured their inter-dependencies, altering how different sectors of the broader US economy interacted. In the aftermath, with accruing data, researchers were naturally interested in ascertaining the nature of the changes, not only in the individual indicators but also in their interconnections. We investigate the hypothesis of pre- and post-interconnections being the same by studying the partial correlation graph among the economic indicators for different market sectors.
In periods between successive extreme events, the economic system usually follows a more predictable set of dynamics that can be modeled as part of a stationary process (after suitable transformations). To analyze an event’s impact, one can use independent models to represent the stationary dynamics before and after the event and compare the parameters of the two models within a Bayesian framework. A change point model is often used to determine structural changes in the time series dynamics; Bai et al. (2023). However, for a significant and catastrophic event like the Great Recession or the COVID-19 pandemic, one typically requires multiple change points to adequately capture the effect of all the shocks to the system. The epochs for the multiple change points are generally unknown. During the Great Recession, the regulatory agencies took several actions whose dates and times are known. However, several additional shocks to the system contributed to the downturn. Accurate testing of hypotheses on structural changes could be challenging without knowing the number and location of change points. Moreover, the period during the event will have very different dynamics from the stable dynamics to which the driving indicators are expected to return after the event. The advantage of a Bayesian paradigm is that we can effectively draw inferences on the impact by studying the posterior of the differences in model parameters from the before and after periods, thereby identifying and quantifying the changes induced by the disruptive event. This approach allows one to quantify the shift in the economic system and glean insight into how the event has changed the underlying dynamics of the system.
We use the FRED-QD data (McCracken and Ng, 2020) for our investigation. The data comprise 248 different economic time series observed at a quarterly frequency. The FRED-QD time series has often been modeled using vector autoregression or its variants McCracken and Ng (2020); Chan (2023); Samadi and Herath (2024). We use the vector autoregression to model the data and investigate the change in the interconnections among the different component series within the VAR framework. Specifically, we study the change in the stationary precision matrix due to the great recession period (December 2007–-June 2009) in a graphical model set up. The graphical model is ideal for studying the interconnection among the economic indicators and any potential changes that may have occurred post-recession. While a change point analysis in vector autoregression can test hypotheses of structural changes, for events spanning a long time window including multiple time points, such as the Great Recession, the specification of a point null hypothesis may be tenuous. Moreover, assumptions such as stability and stationarity are important properties to analyze under a VAR framework (Wang et al., 2022; Zheng, 2024). Thus, we focus on investigating the difference in model parameters between the pre- and the post-recession periods, where the assumption of stationarity (after transformation) may be more meaningful. Given the duration of the crisis, the periods before and after the recession can be reasonably assumed to be independent for VAR modeling purposes. We assume that the underlying dynamics of the economy have recovered to a stable state in the period following the recession, possibly to a different stable state compared to before the financial crisis. We compare independent causal VAR models from the two periods in a Bayesian framework and make inference on the changes in the stationary graphs given by the corresponding stationary precision matrices. The joint analysis of the before and after states is somewhat reminiscent of usual difference-in-difference analysis Angrist and Pischke (2009) done to evaluate the impact of an extreme shock to a dynamical system.
@@
Graphical models are popular models that encode scientific linkages between variables of interest through a conditional independence structure and provide a parsimonious representation of the joint probability distribution of the variables, mainly when the number of variables is large. Thus, learning the graph structure underlying the joint probability distribution of an extensive collection of variables is an important problem, and it has a long history. When the joint distribution is Gaussian, learning graphical structure can be done by estimating the precision matrix leading to the popular Gaussian Graphical Model (GGM); even in the non-Gaussian case one can think of encoding the graph using the partial correlation structure and learn the Partial Correlation Graphical Model (PCGM) by estimating the precision matrix. However, estimating the precision matrix efficiently from the sample can be challenging when the sample has temporal dependence. If estimating the graphical structure is the main objective, the temporal dependence can be treated as a nuisance feature and ignored in graph estimation. When the temporal features are also important to learn, estimating the graph structure and the temporal dependence simultaneously can be considerably more complex.
@@
There are several formulations of graphical models for time series. A time series graph could be one with the nodes representing the entire coordinate processes. In particular, if the coordinate processes are jointly stationary, this leads to a ‘stationary graphical’ structure. Conditional independence in such a graph can be expressed equivalently in terms of the absence of partial spectral coherence between two nodal series; see Dahlhaus (2000). Estimation of such stationary graphs has been investigated by Jung et al. (2015), Basu et al. (2015), Ma and Michailidis (2016), Fiecas et al. (2019) and Basu and Rao (2023).
@@
A weaker form of conditional coding for stationary multivariate time series is a ‘contemporaneous stationary graphical’ structure, where the graphical structure is encoded in the marginal precision matrix. In a contemporaneous stationary graph, the nodes are the coordinate variables of the vector time series. If the time series is Gaussian, this leads to contemporaneous conditional independence among the coordinates of the vector time series. This is the GGM structure, which has the appealing property that the conditional independence of a pair of variables given the others is equivalent to the zero value of the corresponding off-diagonal entry of the precision matrix. Hence, graph learning can be achieved via the estimation of precision matrices. A rich literature on estimating the graphical structure can be found in classical books like Lauritzen (1996); Koller and Friedman (2009); see also Maathuis et al. (2019) and references therein for more recent results. When the distribution is specified only up to the second moment, common in the study of second-order stationary time series, one could learn the contemporaneous partial correlation graphical structure by estimating the stationary precision matrix. In this paper, we use the contemporaneous precision matrix of a stationary time series to estimate the graph along the line of Qiu et al. (2016); Zhang and Wu (2017). We use a vector autoregression (VAR) model to model the temporal dynamics.
@@
To study the graphical relations and changes within the macroeconomic series, this paper combines two popular models: the PCGM for graph estimation and the VAR model for estimating the temporal dynamics in vector time series. In the context of our specific investigation, we require that the processes over the study periods are stable, as encoded by the causality of the VAR process. Also, given the high dimensionality of the problem, dimension reduction in the form of reduced rank VAR with sparse precision matrix is desired. Modeling the partial correlation graph and the temporal dependence with the VAR structure simultaneously is challenging, particularly under constraints such as reduced rank and causality on the VAR model and sparsity on the graphical model. This is achieved in this paper via a novel parameterization of the VAR process. The main contribution of the paper is a methodology that allows for meeting the following two challenging objectives simultaneously:
@@
    Estimating the contemporaneous stationary graph structure under a sparsity constraint.
@@
    Estimating VAR processes with a reduced rank structure under causality constraints.
@@
    establish posterior concentration under priors based on the new parameterization.
2 FRED-QD data and Exploratory Analysis
We study the change in the economic output due to shocks from the Great Recession period (December 2007-June 2009) using the FRED-QD data (McCracken and Ng, 2020). This dataset contains a total of 248 quarterly frequency series that are further divided into 14 disjoint groups, namely 1) NIPA (23 variables); 2) Industrial Production (16 variables); 3) Employment and Unemployment (50 variables); 4) Housing (13 variables); 5) Inventories, Orders, and Sales (9 variables); 6) Prices (48 variables); 7) Earnings and Productivity (14 variables); 8) Interest Rates (19 variables); 9) Money and Credit (14 variables); 10) Household Balance Sheets (9 variables); 11) Exchange Rates (4 variables); 12) Other (2 variables); 13) Stock Markets (3 variables); and 14) Non-Household Balance Sheets (13 variables). The numbers in parentheses indicate the variable counts after excluding those with missingness. The group ‘other’ only contains two variables and is thus excluded from the analysis. The Unemployment and the Prices groups have close to 50 series. Since the number of time points is limited and the groups can be naturally split into further economic subgroups, we decided to split the Unemployment and the Price groups into two parts each. The parts of the Unemployment group comprise all series related to Employment volume (22 variables) (Employment) and those related to unemployment and workforce participation (Unemployment and Participation). In contrast, the parts of the Price group comprise series related to Personal Consumption and Expenditure, with 20 variables (Prices(PCE)), and the prices series related to Consumer Price Indices, with the remaining 28 series (Prices(CPI)). Thus, in our analysis, we have 15 different groups. While estimating the post-recession parameter, we extract the data for 43 time points from July 2009 to January 2020, avoiding the start of the COVID-19 pandemic. We extract the data for 43 time points preceding December 2007 for the pre-recession parameter estimation to have comparable lengths. As a pre-processing step, we use transformations for each 248 time series as outlined in McCracken and Ng (2020).
As an exploratory analysis, we look at the sample estimates of the stationary precision matrices for the time series within each group and each period. Specifically, we perform an informal yet illustrative comparison of each economic group’s precision matrices from before and after the recession. Figure 1 shows the empirical precision estimates for the groups NIPA, Employment, Housing, Prices (CPI), and Money and Credits from the pre- and post-recession periods. The images for the remaining groups are given in Section 3 of the supplement. From the estimates, it is clear that the great recession’s impact was substantial across most of the economic sectors. It is expected to see pervasive changes in the interdependence among the economic indicators since the Great Recession is known to have affected every aspect of the economy Weinberg (2013). This is apparent from the estimated precision matrices. Some sectors are impacted more than others, and the interdependencies among the indicators show more resilience in some sectors, such as exchange rates. The exchange rates showed some degree of resistance; Fenn et al. (2009) demonstrates that certain exchange rate interdependencies remained stable, while others evolved significantly during the credit crisis, highlighting the resilience and adaptability of specific currency relationships. These facts are borne out somewhat in a visual inspection of the empirical estimates of the precision matrices from periods before and after the recession. However, a more rigorous analysis is needed to substantiate or refute such hypotheses.
Figure 1: Estimated sample precisions for some example groups
We use existing modeling tools to perform a model-based analysis to estimate changes in these group-specific time series before and after the recession period. There is evidence that vector autoregressive models, after suitable transformations, adequately describe stationary dynamics in most macroeconomic series Bańbura et al. (2010). Given that some of the groups have more than twenty series, and that the period of the series is limited, VAR modeling needs to be augmented with some dimension-reduction technique before the entire group can be modeled jointly. To this end, we apply the Sparse-VAR(1) model to the before and after data fragments using the R package sparsevar (Vazzoler, 2021) with the default elastic net penalty. To illustrate how the model fits from the pre- and post-recession periods differ, we look at the change in the estimated marginal precision, computed based on the estimated VAR parameters. To this end, for each FRED-QD group $g$, we compute $\hat{\bm{\Omega}}_{g,b}$, the precision matrix from before the recession, and $\hat{\bm{\Omega}}_{g,a}$, the precision matrix from after the recession period. We then compute the precision difference matrix ${\hat{\bm{\Omega}}}^{d}_{g}={\hat{\bm{\Omega}}}_{g,a}-{\hat{\bm{\Omega}}}_{g,b}$. We then define a normalized edge difference value by dividing the off-diagonal entries of the precision difference matrix by the square root of the associated diagonal entries of the sum of the precision matrices, ${\hat{\bm{\Omega}}}^{s}_{g}={\hat{\bm{\Omega}}}_{g,b}+{\hat{\bm{\Omega}}}_{g,a}$. Specifically, we define the scale-free edge-difference parameter as
$\theta_{g}(i,j)={\hat{\Omega}}^{d}_{g,i,j}/[{\hat{\Omega}}^{s}_{g,i,i}{\hat{\Omega}}^{s}_{g,j,j}]^{1/2}$ (2.1)
where for a matrix $\bm{A}$, $A_{i,j}$ denotes the $(i,j)$th entry of $\bm{A}$. Subsequently, we report the percentage of edge-difference parameters $\theta_{g,i,j}$ that were bigger than a threshold $\tau$ in absolute value. For the present analysis, we choose $\tau=0.1$. The aggregate percentage of significant edge-difference values for each group is reported in Table 1. Although the reported change scores depend on the choices of $\tau$, the relative rankings of the groups from high to low changes remained stable even for other choices.
Table 1: Change from pre to post-recession for different economic groups based on VAR(1) estimates
Change-score
National income and product accounts 0.05
Industrial Production 0.22
Employment 0.06
Unemployment and Participation 0.10
Housing 0.21
Inventories, Orders, and Sales 0.33
Prices (PCE) 0.06
Prices (CPI) 0.07
Earnings and Productivity 0.04
Interest Rates 0.13
Money and Credit 0.10
Household Balance Sheets 0.06
Exchange Rates 0.17
Stock Markets 0.33
Non-Household Balance Sheets 0.23
Some groups show appreciable differences from pre- to post-recession, and these findings align with the visual inspection of the empirical precision estimates. However, contrary to the visual findings, several groups do not show significant differences between the pre- and post-recession SparseVar output. These are false negatives, as there is a body of economic research pointing to possibly structural changes in the interdependencies from pre- to post-recession in these component series. The SparseVar findings contradict what is known from economic theory and show the limitations of approaches for estimating changes in interdependencies among the variables when sample dependence is present. The relationship among indicators in Housing is known to have been severely disrupted by the recession (mainly since the initial dynamics of the recessive economy are attributed to the collapse in the housing market), taking several years in the post-recession period to recover, but possibly at a different level. While the sparse VAR does detect some changes in the interdependencies in the housing market, the degree of change is severely underestimated. This highlights the need for a more flexible method for graph estimation in the presence of VAR-type dependence. Furthermore, the sparse VAR estimation procedure sparsevar does not necessarily produce stationary estimates even when the working hypothesis is that the series, after the transformations recommended in McCracken and Ng (2020), are stable in the respective periods. This is another deficiency that is overcome by the proposed model.
@@ 3 Partial Correlation Graph Under Autoregression
In many contexts, the set of variables of interest evolves and is temporally dependent. If the process $\{\bm{X}_{t}:t=0,1,2,\ldots\}$ is stationary, that is, the joint distributions remain invariant under a time-shift, the relational graph of $\bm{X}_{t}$ remains time-invariant. A vector autoregressive (VAR) process provides a simple, interpretable mechanism for temporal dependence by representing the process as a fixed linear combination of itself at a few immediate time points plus an independent random error. It is widely used in time series modeling. In this paper, we propose a Bayesian method for learning the relational graph of a stationary VAR process.
@@
Causality is a property that plays an important role in multivariate time series models, particularly in terms of forecasting. For a causal time series, the prediction formula includes current and past innovations, and hence, a causal time series allows a stable forecast of the future in terms of present and past data. However, the condition of causality imposes complex constraints on the parameters of the process, often making it extremely difficult to impose causality during the estimation process. A practical approach is to parameterize the constrained parameter space of causal processes in terms of unconstrained parameters and write the likelihood and the prior distribution in terms of the unconstrained parameters.
@@
$\det(\bm{A}(z))=0\mbox{ where }\bm{A}(z)=\bm{I}-\bm{A}_{1}z-\cdots-\bm{A}_{p}z^{p},\;z\in{\mathbb{C}},$ (3.2)
the matrix polynomial associated with the VAR equation. For the VAR model to be causal, all roots of the determinantal equation must lie outside the unit disc $\mathcal{D}=\{z\in{\mathbb{C}}:|z|\leq 1\}$. When the roots of the determinantal equation lie outside the unit disc, the associated monic matrix polynomial ${\tilde{\bm{A}}}(z)=z^{p}-\bm{A}_{1}z^{p-1}-\cdots-\bm{A}_{p}$ is called Schur stable. Roy et al. (2019) provided a parameterization of the entire class of Schur stable polynomials and thereby parameterized the space of causal VAR models.
@@
Our main objectives in this paper are to estimate the stationary precision matrix of the process, $[{\rm Var}(X_{t})]^{-1}=\Omega=\bm{\Gamma}^{-1}(0)$, under sparsity restrictions. In the Roy et al. (2019) parameterization, the stationary variance $\bm{\Gamma}(0)$ and hence $\bm{\Omega}$ are functions of the basic free parameters and hence not suitable for estimation of the graphical structure with desired sparsity properties. We suggest a novel modification of the previous parameterization that achieves the goal of parameterizing the graphical structure and the VAR correlation structure directly and separately, thereby facilitating the estimation of both components under the desired restrictions, even in higher dimensions.
@@ 4 Reduced Rank Parameterization and Priors
The main idea in the proposed parameterization is the separation of the parameters pertaining to the graphical structure, $\bm{\Omega},$ and the parameters used to describe the temporal correlation present in the sample. The following result is essential in developing the parameterization. It follows easily from (3.5), but we state it formally due to its central nature in the new parameterization.
@@
Constraint-free parameterization: Based on Proposition 1, for a causal VAR process the successive differences $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ are nonnegative definite for $j=1,\ldots,p$, where $\bm{\Omega}=\bm{C}_{0}^{-1}$. Hence, these successive differences and the precision matrix $\bm{\Omega}$ can be parameterized using an unrestricted parameterization that maps non-negative definite matrices to free parameters.